OpenStudy (castiel):

sin^4(x)+cos^4(x) = 5/8

OpenStudy (castiel):

how to solve this?

OpenStudy (anonymous):

mathway.com

OpenStudy (anonymous):



OpenStudy (anonymous):

Let cos^2(x)=a Sin^2(x)=1-cos^2(x)=1-a Then we have (sin^2(x))^2+a^2=5/8 This is. (1-a)^2+a^2=5/8 I trust you can do the rest